Activity Rates and Activity-Based Product Costing

Hammer Company produces a variety of electronic equipment. One of its plants produces two laser printers: the deluxe and the regular. At the beginning of the year, the following data were prepared for this plant:

Deluxe Regular
Quantity 100,000 800,000
Selling price $900 $750
Unit prime cost $529 $483

In addition, the following information was provided so that overhead costs could be assigned to each product:

Activity Name Activity Driver Deluxe Regular Activity Cost
Setups Number of setups 300 200 $1,850,000
Machining Machine hours 100,000 300,000 40,000,000
Engineering Engineering hours 50,000 100,000 13,500,000
Packing Packing orders 100,000 400,000 450,000

Required:

1. Calculate the overhead rates for each activity. If required, carry your answers out to the nearest cent.

Setups $ per setup
Machining $ per machine hour
Engineering $ per engineering hour
Packing $ per packing order

2. Calculate the per-unit product cost for each product. Round your answers to the nearest whole dollar.

Deluxe $per unit
Regular $per unit
Answer #1

Activity rate = overhead cost/total activity

Setups = 1850,000/(300+200) =$3700 per set up

Machining = 40,000,000/(100,000+300,000) =$10 per machine hour

Engineering = 13500,000/150,000 =$90 per hour

Packing = 450,000/500,000 =$0.9 per order

2. Deluxe

Unit cost = 529+ (300*3700+100,000*10+50,000*90+100,000*0.9)/100,000

=$596 per unit

Regular = 499.375